Last 10 Maintenance posts, sorted by last updated
Order posts by limited to posts

MAINTENANCE Completed TalkTalk
AFFECTING
TalkTalk
STARTED
Jul 11, 11:11 AM (3½ days ago)
CLOSED
Jul 11, 11:50 AM (3½ days ago)
DESCRIPTION
At 11:21 we saw a small number of TalkTalk lines drop and reconnect. We're investigating.
Resolution: A TalkTalk backhaul connection failed (twice?) and traffic was re-routed, as a result a small number to TalkTalk connections dropped and reconnected.

MAINTENANCE Completed VoIP and SIMs
AFFECTING
VoIP and SIMs
STARTED
Jul 08, 10:30 AM (6½ days ago)
CLOSED
Jul 08, 05:05 PM (6¼ days ago)
DESCRIPTION
ONSIM are making changes to address some long standing issues with slow connects and multiple connects for outgoing calls. Sorry for the short notice, they have just advised us. We have been nagging them to fix this for some time. It is possible some calls from mobiles may fail, but it is worth retrying if they do. We are hoping this will only last a few minutes, but not depends on the mobile carrier completing the work promptly.
Resolution: We have not had feedback, but tests suggest calls are being consistently reliable now.

MAINTENANCE Completed Broadband
AFFECTING
Broadband
STARTED
Jun 14, 10:00 AM (30½ days ago)
CLOSED
Jul 08, 10:45 AM (6½ days ago)
DESCRIPTION

As you are probably aware A&A develops its own network hardware under the FireBrick brand. We have been, for some time now, using a mixture of the older "FB6000" and the newer "FB9000" models as BGP routers, L2TP routers and LNSs which terminate customer broadband connections; For the most part the FB9000 LNSs have been is use for customers with faster speed lines. We've been using FB9000s for our L2TP services and as BGP routers for over a year now.

You might have seen that we did have some stability problems earlier in the roll out of the FB9000 for our LNS platform and therefore paused the deployment. For some months now, we have running firmware with well established stability and as such will be rolling out FB9000s in replacement of all older FB6000 LNS. This work will happen in small stages during June and July.

This will be mildly customer impacting, causing a brief outage as each line moves across, but no worse than any common LNS firmware update like we have been doing consistently for years. We will do migrations in the early hours of the morning and according to your preferences in the control pages.

We will post updates to this status post when parts of this wider project are scheduled. eg when we schedule moving a set of customers off a specific LNS on to a new one.


Resolution: These migrations have completed. All broadband customers are now on FB9000 LNSs.

MAINTENANCE Completed Control Pages
AFFECTING
Control Pages
STARTED
Jun 04, 08:00 PM (1¼ months ago)
CLOSED
Jun 04, 08:26 PM (1¼ months ago)
DESCRIPTION
On Tuesday evening we'll be doing some maintenance on the Control Pages which will mean it will be our of action for for a little while as this work is carried out.
Resolution:

MAINTENANCE Completed BT
AFFECTING
BT
STARTED
May 24, 12:01 AM (1¾ months ago)
CLOSED
May 24, 01:00 AM (1¾ months ago)
DESCRIPTION
BT have planned work on their side of one of our hostlinks, on 24th May between midnight and 6AM. We will move traffic away from this hostlink beforehand so as to minimise the impact on customers. We don't expect this to impact customers.
Resolution: The work was carried out without affecting customer traffic.

MAINTENANCE Completed VoIP and SIMs
AFFECTING
VoIP and SIMs
STARTED
May 01, 06:00 PM (2¼ months ago)
CLOSED
May 06, 05:00 PM (2¼ months ago)
DESCRIPTION
Following issues this week with calls to SIP2SIM phones not working reliably, the carrier has finally updated their systems to have fixed firewall rules. This solves one of the key issues we have seen with unreliability. Their firewall rules were not properly made from the DNS records for some reason. We are now able to finish updating our routing logic over the next few days. This should not impact services, but there is a risk. Thankfully the risk is quite low, and importantly likely short lived, as it seems their main call handling servers are reasonably quick to pick up DNS changes (unlike their firewall rules). The aim is to make most of the changes over the weekend.
Resolution:

MAINTENANCE Completed TalkTalk
AFFECTING
TalkTalk
STARTED
Apr 21, 09:00 PM (2¾ months ago)
CLOSED
May 16, 06:00 AM (2 months ago)
DESCRIPTION

We have multiple interlinks to TalkTalk that carry our broadband traffic. TalkTalk have scheduled planned work on both these links during a four week period from Tuesday 23rd April until 16th May. (Specifically midnight to 6AM on 23rd, 25th, 25th April and 1st, 2nd, 9th, 16th May.

Due to the work being carried out (Software updates of their "LTSs") we are unable to move traffic seamlessly between our interlinks and so TalkTalk customers will see their connections drop and reconnect on these early mornings.


Resolution: TalkTalk's PEW window is now over.

MAINTENANCE Completed CityFibre
AFFECTING
CityFibre
STARTED
Apr 20, 04:00 AM (2¾ months ago)
CLOSED
Apr 22, 07:25 AM (2¾ months ago)
DESCRIPTION
We'll be performing some work on one of our routers for CityFibre connections. Some CityFibre customers will see their connection drop and reconnect moments after at 4AM on Saturday morning.
Resolution: This work has been completed.

MAINTENANCE Completed Ethernet BGP
AFFECTING
Ethernet BGP
STARTED
Apr 15, 10:30 PM (3 months ago)
CLOSED
Apr 17, 09:30 PM (2¾ months ago)
DESCRIPTION
We're performing some maintenance on the primary router used for our Ethernet (Etherway/Etherflow) services and our customer BGP sessions (eg customers with their own BGP sessions to us) - "A.Weightless". We have moved traffic on to our secondary router for the next 48 hours. This move is seamless and customer traffic is not affected.
Resolution: This has been completed.

MAINTENANCE Completed L2TP
AFFECTING
L2TP
STARTED
Apr 10, 04:00 AM (3 months ago)
CLOSED
Apr 11, 04:10 AM (3 months ago)
DESCRIPTION
We will be replacing the hardware of our main L2TP router during the day on Wednesday 10th April. As part of this work we will be moving L2TP customers over to the backup L2TP server shortly after 4AM on 10th April. This was cause customers to drop and reconnect.
Resolution: This has been completed.